OpenStudy (anonymous):

P is proportional to q^3 It is known that p=24 for a particular value of q. Find the value of p wwhen q is halved

OpenStudy (yttrium):

\[p = kq^3\] \[24 = kq^3\] \[24 = k (q/2)^3\] \[24 = k(q^3/8)\] Therefore, we can conclude that \[p = 2/8 = \] If you have more questions you can pm me.

OpenStudy (yttrium):

i mean p = 24/8 = 3